model to pedia ere quali, measured on a scale tom O (very bad) to 10 (ascenont), based on aloohol cortent0%). Corlete pats(a)trough, (d) below. O A. The siope, by, implies that the alcohol content is equal to the value ofb, in percentages D. The ^□(Round to two decimal places as needed )
Answer #1

c) The regression equation is \hat{y1} = b0 + b1*x

x = 11%

b0 = 5.1

b1 = 0.2

\hat{y1} = 5.1 + 0.2*11 = 7.3

If we take x in the form of decimal we get x = 0.11

then value of \hat{y1} will be change.

\hat{y1} = 5.1 + 0.2*0.11 = 5.12